OpenStudy (anonymous):

find both coordinates of the vertex of thenparabola g(x)=-2x^2+2x-5

OpenStudy (osanseviero):

Having a function F(x) = ax^2 + bx + c You can find the parabola: V ( -B/2A , F(-B/2A))

OpenStudy (solomonzelman):

re-type the question.

OpenStudy (anonymous):

g(x)=-2x^2+2x-5

OpenStudy (rizags):

find the axis of symmetry first

OpenStudy (anonymous):

how

OpenStudy (rizags):

well, for a parabola in the form ax^2+bx+c, it is x=-b/2a

OpenStudy (anonymous):

thankyou

OpenStudy (rizags):

what did you get?

OpenStudy (anonymous):

1/2

OpenStudy (jdoe0001):

\(\bf \textit{vertex of a parabola}\\ \quad \\ g(x) = {\color{red}{ -2}}x^2{\color{blue}{ +2}}x{\color{green}{ -5}}\qquad \left(-\cfrac{{\color{blue}{ b}}}{2{\color{red}{ a}}}\quad ,\quad {\color{green}{ c}}-\cfrac{{\color{blue}{ b}}^2}{4{\color{red}{ a}}}\right)\)

OpenStudy (rizags):

ok, now plug that in for x, into the equation and tell me what you get

OpenStudy (jdoe0001):

or do as Rizags suggested

OpenStudy (anonymous):

-9/2

OpenStudy (rizags):

yes, correct, now the vertex will be (x, g(x)), or \[(\frac{1}{2}, -\frac{9}{2})\]

OpenStudy (anonymous):

thank you so much

OpenStudy (rizags):

yup, no prob